# Jasmine: Harnessing Diffusion Prior for Self-Supervised Depth Estimation

### v2

@@ 1 Introduction
Estimating depth from monocular images is a fundamental problem in computer vision, which plays an essential role in various downstream applications such as 3D/4D reconstructiondepthsplat , autonomous drivingm4depth , etc. Compared with supervised methodsmarigold ; e2eft ; lotus , self-supervised monocular depth estimation (SSMDE) mines 3D information solely from video sequences, significantly reducing reliance on expensive ground-truth depth annotations. These methods derive supervision from geometric constraints (e.g., scene depth consistency) through cross-frame reprojection loss, and the ubiquitous video data further suggests an unlimited working potential. However, view reconstruction-based losses suffer from occlusions, texture-less regions, and illumination changesmonovit , which severely restrict the model’s capacity to recover fine-grained details and may cause pathological overfitting to specific datasets.
@@ 2 Relative Work
As the diffusion paradigm showcases its talents in generative tasksddpm ; ddim ; liao2024denoising ; yue2024arbitrary , DDPDDP first reformulates depth perception as a depth map denoising task and leads to giant progress. Followers like DDVMddvm , MonoDiffusionmonodfs , and D4RDd4rd (the latter two are self-supervised methods but employ self-designed diffusion) all demonstrate the advantages of this paradigm in various MDE sub-tasks. Subsequently, the most renowned diffusion model, Stable Diffusionsd , has demonstrated significant potential for depth perception tasks. VPDvpd , TAPDtapd , and Prior-Diffusionpdfs use Stable Diffusion as a multi-modal feature extractor, leveraging textual modality information to improve depth estimation accuracy. Concurrently, Marigoldmarigold and GeoWizardgeo enhanced model generalization and detail preservation by fine-tuning Stable Diffusion, capitalizing on its prior training with large-scale, high-quality datasets. Afterward, E2E FTe2eft and Lotuslotus further accelerated inference by optimizing the noise scheduling process. In this work, our Jasmine continues these works and extends SD to the field of self-supervision.
@@ Appendix A Proof of Scale-Invariant Depth
Note that for transformations within the same scene, the relative pose R remains constantssi2si . Also, $a$ and $b$ here can be arbitrary numbers, so we have:
@@
ManyDepthmanydepth CVPR2021 ( -1,0) K(40K+0) 0.091 0.694 4.245 0.171 0.911 0.968 0.983
Dual Refinedualrefine CVPR2023 (-1,0) K(40K+0) 0.087 0.674 4.130 0.167 0.915 0.969 0.984
Mono-ViFIvifi ECCV2024 (-1,0,1) K(40K+0) 0.089 0.556 3.981 0.164 0.914 0.971 0.986
EPCDepthepcdepth ICCV2021 Stereo K(45K+0) 0.091 0.646 4.207 0.176 0.901 0.966 0.983
PlaneDepthplanedepth CVPR2023 Stereo K(45K+0) 0.085 0.563 4.023 0.171 0.910 0.968 0.984
@@ B.1 Step-by-Step Workflow
  • Step 3: Self-Supervised Signal Generation (Image Reprojection):
@@ C.1 SGD Loss for Dynamic Object
Building on this, we adopt the GDS loss (only the first stage) fromgds to handle dynamic objects. This approach is based on the smoothness loss and introduces a ground-contact-prior mask $M_{gr}$, defined as:
@@ C.4 Eliminating the Bottleneck of Teacher Loss
Furthermore, to avoid the performance ceiling imposed by the teacher model, we draw inspiration from D4RDd4rd , and employ an adaptive filtering mechanism:
@@
The segmentation model mentioned above is Mask2Formermask2former throughout. To avoid the misconception that the details stem from the segmentation network’s output, we further expand Fig. 3 (in paper) with segment-based methods. In fact, to preserve these details, a naive approach is to employ semantic segmentation constraintsgds . However, as shown in Fig. 8 (a-2), using the semantic triplet loss from gds not only disrupts the depth distribution but also introduces spurious edges, rendering it incompatible with SD’s latent priors. Furthermore, the results in paper Table 1 are compared with a segmentation-based model Jaeon et al*, and our method achieves superior performance.
@@ Appendix D Preliminaries of GRU
The Gated Recurrent Unit (GRU) is a type of recurrent neural network (RNN) that uses gating mechanisms to control the flow of information between the previous hidden state and the current input, making it effective for sequence modeling tasks. Recently, GRUs have been gradually used in depth estimationiebins ; nddepth . A GRU cell updates its hidden state $\mathbf{h}^{k+1}$ based on the previous hidden state $\mathbf{h}^{k}$ and the current input $\mathbf{x}^{k}$. This update process is primarily managed by two gates: the reset gate ($\mathbf{r}$) and the update gate ($\mathbf{z}$). The reset gate determines how much of the past information (from the previous hidden state) to effectively "forget" or "reset" before computing a new candidate hidden state. The update gate then decides how much of this new candidate hidden state should be incorporated into the final hidden state, versus how much of the previous hidden state to carry over. The mathematical formulation for these operations is as follows:
$\begin{array}[]{l}\mathbf{z}^{k+1}=\sigma([\mathbf{h}^{k},\mathbf{x}^{k}],W_{z}),\quad\mathbf{r}^{k+1}=\sigma([\mathbf{h}^{k},\mathbf{x}^{k}],W_{r}),\\ \widehat{\mathbf{h}}^{k+1}=\tanh({Conv}([\mathbf{r}^{k+1}\odot\mathbf{h}^{k},\mathbf{x}^{k}],W_{h})),\\ \mathbf{h}^{k+1}=(1-\mathbf{z}^{k+1})\odot\mathbf{h}^{k}+\mathbf{z}^{k+1}\odot\widehat{\mathbf{h}}^{k+1},\end{array}$ (23)
@@ E.1 Evaluation on KITTI Improved Benchmark
The standard self-supervised evaluation on the KITTI dataset is typically conducted using the raw LiDAR GT. However, due to the noisy nature of LiDAR data and known preprocessing issues [^1] Discussed in MonoDepth2 repository: https://github.com/nianticlabs/monodepth2/issues/274, we also provide the eigen improved benchmark result. Following the practice adopted in Mono-ViFi vifi , as shown in Table 5, this evaluation further confirms Jasmine’s SoTA performance, where it consistently outperforms other leading methods.
Table 5: Quantitative results on the KITTI dataset using the improved GT impr . The performance of SD-based methods is not fully reported in their original papers.
@@
Marigold marigold 0.099 - - - 0.916 0.987 -
E2E-FT e2eft 0.096 - - - 0.921 0.980 -
Lotus lotus 0.081 - - - 0.931 0.987 -
@@
MonoViT monovit 0.068 0.314 3.125 0.105 0.948 0.992 0.998
MonoViFi vifi 0.071 0.338 3.539 0.113 0.937 0.990 0.998
@@ Appendix G Evaluation Metrics
Similar to mdp2 , we employ the following evaluation metrics in our experiments,
@@ Appendix I Limitation and Future Work
As noted in Sec. 2, the ubiquitously available videos suggest that the self-supervised methods possess significant data advantages and working potential. In this paper, Jasmine is trained on only tens of thousands of data samples and only on a driving dataset (KITTI), leaving room for further exploration in scaling up training data to other domains (industry, indoor, etc). We believe that if there really emerges a ‘GPT’ moment for 3D perception in the future, it will more likely involve self-supervised methods trained on videos rather than learning from annotated depth. Furthermore, we believe that the unsupervised Stable Diffusion fine-tuning paradigm proposed in this paper can be applied to other related fields, such as depth completion yan1 ; yan2 ; yan3 , depth super-resolution yan4 ; yan5 , and Multi-view Stereo yuan1 ; yuan2 ; yuan3 ; yuan4 .

### v3

@@ 1 Introduction
Estimating depth from monocular images is a fundamental problem in computer vision, which plays an essential role in various downstream applications such as 3D/4D reconstructiondepthsplat ; xgk2 , autonomous drivingm4depth , etc. Compared with supervised methodsmarigold ; e2eft ; lotus , self-supervised monocular depth estimation (SSMDE) mines 3D information solely from video sequences, significantly reducing reliance on expensive ground-truth depth annotations. These methods derive supervision from geometric constraints (e.g., scene depth consistency) through cross-frame reprojection loss, and the ubiquitous video data further suggests an unlimited working potential. However, view reconstruction-based losses suffer from occlusions, texture-less regions, and illumination changesmonovit , which severely restrict the model’s capacity to recover fine-grained details and may cause pathological overfitting to specific datasets.
@@ 2 Relative Work
As the diffusion paradigm showcases its talents in generative tasksddpm ; ddim ; liao2024denoising ; yue2024arbitrary ; wang0 ; xgk1 , DDPDDP first reformulates depth perception as a depth map denoising task and leads to giant progress. Followers like DDVMddvm , MonoDiffusionmonodfs , and D4RDd4rd (the latter two are self-supervised methods but employ self-designed diffusion) all demonstrate the advantages of this paradigm in various MDE sub-tasks. Subsequently, the most renowned diffusion model, Stable Diffusionsd , has demonstrated significant potential for depth perception tasks. VPDvpd , TAPDtapd , and Prior-Diffusionpdfs use Stable Diffusion as a multi-modal feature extractor, leveraging textual modality information to improve depth estimation accuracy. Concurrently, Marigoldmarigold and GeoWizardgeo enhanced model generalization and detail preservation by fine-tuning Stable Diffusion, capitalizing on its prior training with large-scale, high-quality datasets. Afterward, E2E FTe2eft and Lotuslotus further accelerated inference by optimizing the noise scheduling process. In this work, our Jasmine continues these works and extends SD to the field of self-supervision.
@@ Appendix A Proof of Scale-Invariant Depth
Note that for transformations within the same scene, the relative pose R remains constant[18]. Also, $a$ and $b$ here can be arbitrary numbers, so we have:
@@
ManyDepth[55] CVPR2021 ( -1,0) K(40K+0) 0.091 0.694 4.245 0.171 0.911 0.968 0.983
Dual Refine[1] CVPR2023 (-1,0) K(40K+0) 0.087 0.674 4.130 0.167 0.915 0.969 0.984
Mono-ViFI[29] ECCV2024 (-1,0,1) K(40K+0) 0.089 0.556 3.981 0.164 0.914 0.971 0.986
EPCDepth[36] ICCV2021 Stereo K(45K+0) 0.091 0.646 4.207 0.176 0.901 0.966 0.983
PlaneDepth[52] CVPR2023 Stereo K(45K+0) 0.085 0.563 4.023 0.171 0.910 0.968 0.984
@@ B.1 Step-by-Step Workflow
  • Step 3: Self-Supervised Signal Generation (Image Reprojection[48]):
@@ C.1 SGD Loss for Dynamic Object
Building on this, we adopt the GDS loss (only the first stage) from[34] to handle dynamic objects. This approach is based on the smoothness loss and introduces a ground-contact-prior mask $M_{gr}$, defined as:
@@ C.4 Eliminating the Bottleneck of Teacher Loss
Furthermore, to avoid the performance ceiling imposed by the teacher model, we draw inspiration from D4RD[50], and employ an adaptive filtering mechanism:
@@
The segmentation model mentioned above is Mask2Former[6] throughout. To avoid the misconception that the details stem from the segmentation network’s output, we further expand Fig. 3 (in paper) with segment-based methods. In fact, to preserve these details, a naive approach is to employ semantic segmentation constraints[34]. However, as shown in Fig. 8 (a-2), using the semantic triplet loss from [34] not only disrupts the depth distribution but also introduces spurious edges, rendering it incompatible with SD’s latent priors. Furthermore, the results in paper Table 1 are compared with a segmentation-based model Jaeon et al*, and our method achieves superior performance.
@@ Appendix D Preliminaries of GRU
The Gated Recurrent Unit (GRU) is a type of recurrent neural network (RNN) that uses gating mechanisms to control the flow of information between the previous hidden state and the current input, making it effective for sequence modeling tasks. Recently, GRUs have been gradually used in depth estimation[45, 43]. A GRU cell updates its hidden state $\mathbf{h}^{k+1}$ based on the previous hidden state $\mathbf{h}^{k}$ and the current input $\mathbf{x}^{k}$. This update process is primarily managed by two gates: the reset gate ($\mathbf{r}$) and the update gate ($\mathbf{z}$). The reset gate determines how much of the past information (from the previous hidden state) to effectively "forget" or "reset" before computing a new candidate hidden state. The update gate then decides how much of this new candidate hidden state should be incorporated into the final hidden state, versus how much of the previous hidden state to carry over. The mathematical formulation for these operations is as follows:
$\begin{array}[]{l}\mathbf{z}^{k+1}=\sigma([\mathbf{h}^{k},\mathbf{x}^{k}],W_{z}),\qquad\mathbf{r}^{k+1}=\sigma([\mathbf{h}^{k},\mathbf{x}^{k}],W_{r}),\\ \widehat{\mathbf{h}}^{k+1}=\tanh({Conv}([\mathbf{r}^{k+1}\odot\mathbf{h}^{k},\mathbf{x}^{k}],W_{h})),\\ \mathbf{h}^{k+1}=(1-\mathbf{z}^{k+1})\odot\mathbf{h}^{k}+\mathbf{z}^{k+1}\odot\widehat{\mathbf{h}}^{k+1},\end{array}$ (23)
@@ E.1 Evaluation on KITTI Improved Benchmark
The standard self-supervised evaluation on the KITTI dataset is typically conducted using the raw LiDAR GT. However, due to the noisy nature of LiDAR data and known preprocessing issues [^1] Discussed in MonoDepth2 repository: https://github.com/nianticlabs/monodepth2/issues/274, we also provide the eigen improved benchmark result. Following the practice adopted in Mono-ViFi [29], as shown in Table 5, this evaluation further confirms Jasmine’s SoTA performance, where it consistently outperforms other leading methods.
Table 5: Quantitative results on the KITTI dataset using the improved GT [47]. The performance of SD-based methods is not fully reported in their original papers.
@@
Marigold [25] 0.099 - - - 0.916 0.987 -
E2E-FT [33] 0.096 - - - 0.921 0.980 -
Lotus [21] 0.081 - - - 0.931 0.987 -
@@
MonoViT [83] 0.068 0.314 3.125 0.105 0.948 0.992 0.998
MonoViFi [29] 0.071 0.338 3.539 0.113 0.937 0.990 0.998
@@ Appendix G Evaluation Metrics
Similar to [15], we employ the following evaluation metrics in our experiments,
@@ Appendix I Limitation and Future Work
As noted in Sec. 2, the ubiquitously available videos suggest that the self-supervised methods possess significant data advantages and working potential. In this paper, Jasmine is trained on only tens of thousands of data samples and only on a driving dataset (KITTI), leaving room for further exploration in scaling up training data to other domains (industry, indoor, etc). We believe that if there really emerges a ‘GPT’ moment for 3D perception in the future, it will more likely involve self-supervised methods trained on videos rather than learning from annotated depth. Furthermore, we believe that the unsupervised Stable Diffusion fine-tuning paradigm proposed in this paper can be applied to other related fields, such as depth completion [65, 63, 66, 70, 71, 60], depth super-resolution [64, 54], Multi-view Stereo [74, 75, 72, 73], Stereo Matching [56, 58, 57, 53] and Language Aid Depth Estimation [78, 80, 79, 82].